Luthan Company uses a predetermined overhead rate of $23.40 perdirect labor-hour. this predetermined rate was based on 11,000estimated direct labor-hours and $257,400 of estimated totalmanufacturing overhead.

the company incurred actual total manufacturing overhead costs of$249,000 and 10,8000 total direct labor-hours during theperiod.

Determine the amount of manufacturingoverhead that would have been applied to units of product duringthe period.
Answer #1
Concepts and reason

Product cost: This is the total cost incurred to manufacture a product. Product costs include direct materials cost, direct labor costs, and manufacturing (factory) overhead costs.

Direct materials cost: Direct materials cost is the cost ofprime raw materials which are directly related or identified with the goods produced. For example, the cost of plastic component is identified as direct materials cost in the production of plastic bottle.

Direct labor cost: The costs that are related to the labor employed in manufacturing process are known as direct labor costs. The direct labor costs includes not only wages of employees, but also worker’s compensation, life and medical insurance, payroll taxes, training costs, and pension contributions.

Manufacturing overhead costs: The costs, which do not relate directly with the manufacturing of products, are referred to as manufacturing or factory overhead costs or indirect costs. Some examples are insurance cost related to factory operations, depreciation incurred on factory equipment, electricity of the equipment, rent of the plant, and compensation of plant managers.

Fundamentals

Predetermined overhead rate: The cost accountants estimate a rate for all the overhead costs to compute manufacturing overhead. That allocation rate which is computed as a ratio of total estimated overhead costs to total estimated quantity, is known as predetermined overhead rate.The following is the formula to calculate predetermined overhead rate:

Predeterminedoverheadrate}=TotalestimatedoverheadcostsTotalestimatedquantityoftheoverheadallocationbase\left. \begin{array}{l}\\{\rm{Predetermined }}\\\\{\rm{overhead rate}}\\\end{array} \right\}{\rm{ = }}\frac{{{\rm{Total estimated overhead costs}}}}{{{\rm{Total estimated quantity of the overhead allocation base}}}}{\rm{ }}

Manufacturing overhead: The costs which are indirectly related to production process are known as manufacturing costs. The following is the formula to calculate manufacturing overhead allocated:

Manufacturingoverhead}={Predeterminedoverheadrate×Actualquantityofoverheadallocationbase}\left. \begin{array}{l}\\{\rm{Manufacturing }}\\\\{\rm{overhead}}\\\end{array} \right\}{\rm{ = }}\left\{ \begin{array}{l}\\{\rm{Predetermined overhead rate \times }}\\\\{\rm{Actual quantity of overhead allocation base}}\\\end{array} \right\}

In the given problem, the allocation base used to compute predetermined overhead rate is direct labor hours. This could be verified as shown below:

Predeterminedoverheadrate}=TotalestimatedoverheadcostTotalestimateddirectlaborhours=$257,40011,000directlaborhours=$23.40perdirectlaborhour\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Predetermined }}\\\\{\rm{overhead rate}}\\\end{array} \right\}{\rm{ = }}\frac{{{\rm{Total estimated overhead cost}}}}{{{\rm{Total estimated direct labor hours}}}}{\rm{ }}\\\\{\rm{ = }}\frac{{{\rm{\$ 257,400}}}}{{{\rm{11,000 direct labor hours}}}}\\\\{\rm{ = \$ 23}}{\rm{.40 per direct labor hour}}\\\end{array}

Therefore, predetermined overhead rate is $23.40 per direct labor hour.

Compute manufacturing overhead as shown below:

Manufacturingoverhead}={Predeterminedoverheadrate×Actualdirectlaborhours}=$23.40×10,800directlaborhours=$252,720\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Manufacturing }}\\\\{\rm{overhead}}\\\end{array} \right\}{\rm{ = }}\left\{ \begin{array}{l}\\{\rm{Predetermined overhead rate \times }}\\\\{\rm{Actual direct labor hours}}\\\end{array} \right\}\\\\{\rm{ = \$ 23}}{\rm{.40 \times 10,800 direct labor hours}}\\\\{\rm{ = \$ 252,720}}\\\end{array}

Therefore, manufacturing overhead is $252,720.

Ans:

Manufacturing overhead is $252,720.